As a technician in a large pharmaceutical research firm, you need to produce 200. mL of 1.00 mol L−1 potassium phosphate buffer solution of pH = 7.01. The pKa of H2PO4− is 7.21.

You have the following supplies: 2.00 L of 1.00 mol L−1 KH2PO4 stock solution, 1.50 L of 1.00 mol L−1 K2HPO4 stock solution, and a carboy of pure distilled H2O.

How much 1.00 mol L−1 KH2PO4 will you need to make this solution?

Express your answer to three significant digits with the appropriate units.
Answer #1

Ans:

Stock solution-Supply

1.0 M KH2PO4 (2L)

1.0 M K2HPO4 (1.5L)

Required pH of the buffer = 7.01

pKa of (H2PO4)- = 7.21

(H2PO4)- (aq)+ H2O (l) ⇋ (HPO4)2- (aq) + H+ (aq)

pH of this solution if given by Henderson–Hasselbalch equation;

pH = pKa + log([Conjugate base]/[Acid]) = pKa + log([(HPO4)2-]/[(H2PO4)-])

7.01 = 7.21 + log ([(HPO4)2-]/[(H2PO4)-])

[(HPO4)2-]/[(H2PO4)-] = 0.631= [(K2HPO4)]/[(KH2PO4)]

0.631= (No. of moles of K2HPO4/0.2)/ (No. of moles of KH2PO4 /0.2)

No. of moles of K2HPO4 = 0.631 x No. of moles of KH2PO4

Volume of the buffer required = 200 mL = 0.2 L

Volume of 1.0 M KH2PO4 (V1) + Volume of 1.0 M K2HPO4 (V2) = 0.2 L ==> V1 + V2 = 0.2L

We know that volume = Mole/Molarity

V1 = no. of moles of KH2PO4 / 1.0

V2 = no. of moles of K2HPO4 / 1.0 = 0.631 x No. of moles of KH2PO4 / 1.0

Volume of 1.0 M KH2PO4 (V1) + Volume of 1.0 M K2HPO4 (V2) = 0.2 L ==> V1 + V2 = 0.2L

(no. of moles of KH2PO4 / 1.0 ) + (0.631 x No. of moles of KH2PO4 / 1.0) = 0.2

No. of moles of KH2PO4 = 0.2/1.631 = 0.1226 mol

No. of moles of K2HPO4 = 0.631 x No. of moles of KH2PO4 = 0.631 x 0.1226 = 0.0774 mol

V1 = no. of moles of KH2PO4 / 1.0 = 0.1226/1.0 = 0.1226 L

V2 = no. of moles of K2HPO4 / 1.0 = 0.631 x No. of moles of KH2PO4 / 1.0 = 0.0774/1.0 = 0.0774 L

122.6 mL of KH2PO4 (1.0 M) mixed with 77.4 mL of K2HPO4 (1.0 M) gives us 200 mL of buffer with pH = 7.01